# Statistical Collusion by Collectives on Learning Platforms

### v1

@@ 1 Introduction
In a network, agents often have diverse motivations and multiple incentives. When the incentives of the interacting agents do not fully align with those of the designer of the learning system, the former may wish to influence the learning process. This becomes salient in the common case in which a learning system interacts with a large number of agents. Such agents may collaborate, forming a cartel, by pooling their data and devising a common strategy. Accordingly, even if the learning algorithm is robust to single-agent adversarial behavior, the collective may be able to exert a significant influence on the algorithm. This concept of collective action has its origins in economic theories Olson (1965) and has been more recently explored in the context of machine learning Hardt et al. (2023). It is crucial to understand how such collectives can influence learning algorithms to better comprehend the dynamics within the network of agents and to deploy algorithms that are reliable and aligned with the majority interests of consumers. We will focus specifically on the case of a platform that deploys a learning algorithm and interacts with a population of consumers.
@@ 2 Related Work
What sets the concept of collective action apart conceptually is its treatment of the collective as a group of individuals, each representing a data point. This perspective also has economic, social, and political dimensions, as certain groups of individuals can unite and collaborate to influence decisions. Such ideas have been explored in areas of research at the intersection of machine learning and other fields Vincent et al. (2019); Albert et al. (2020); Vincent et al. (2021); Albert et al. (2021); Creager & Zemel (2021); Vincent & Hecht (2021). See Appendix A of Hardt et al. (2023) for an in-depth analysis of the related work on collective action.
An important contribution of Hardt et al. (2023) is to study data poisoning via the formalism of Bayes-optimal classification, which yields a conceptual inversion of the idea of strategic classification Hardt et al. (2016). While strategic classification revolves around a firm’s ability to anticipate and respond to the actions of a single, strategic individual, collective action shifts the focus toward a scenario where individuals collectively anticipate and strategically respond to the optimizing behavior of the firm. This concept has also been explored by Zrnic et al. (2021). Unlike traditional strategic classification, which primarily considers the firm’s perspective, collective action highlights the role of workers and consumers on online platforms.
@@ 3.2.1 Feature-label signal planting
$\begin{split}\hat{S}(n)&\geq\underset{\tilde{x}\sim\tilde{D}^{(n)}}{\hat{\mathbb{P}}}\left[\frac{n}{N}\left(\underset{\tilde{D}^{(n)}}{\hat{\mathbb{P}}}(\tilde{x})-2R_{\tilde{\delta}}(n)\right)-\frac{N-n}{N}\left(\Delta_{\tilde{x}}^{(n)}+2R_{\tilde{\delta}}(n)+2R_{\tilde{\delta}}(N-n)\right)-\frac{\varepsilon}{1-\varepsilon}>0\right]\\ &\quad-R_{\tilde{\delta}}(n)-R_{\tilde{\delta}}(N_{\rm test}),\end{split}$ (1)
@@
Proofs and additional remarks can be found in Appendix E. Note that lower bound (1) is fully computable by the collective as it depends only on datasets $D^{(n)}$ and $\tilde{D}^{(n)}$. $D^{(n)}$ is obtained through data pooling by the collective’s members, and $\tilde{D}^{(n)}$ is computed by applying strategy $h$ to the dataset $D^{(n)}$. Details on the algorithm for computing the lower bound are provided in Appendix B, and the code is available at https://github.com/GauthierE/statistical-collusion.
The interpretation of lower bound (1) is that success increases step by step: each feature $\tilde{x}$ in the signal set $\tilde{\mathcal{X}}$ has a certain resistance to being planted. As the relative size of the collective $n/N$ gradually increases, features $\tilde{x}$ are cracked as their resistance breaks, in decreasing order of resistance. Fo each feature $\tilde{x}$, its resistance depends on three terms. The first one, here $\frac{n}{N}(\underset{\tilde{D}^{(n)}}{\hat{\mathbb{P}}}(\tilde{x})-2R_{\tilde{\delta}}(n))$, represents how prevalent the feature $\tilde{x}$ is in the modified data: the more frequently $\tilde{x}$ appears in the poisoned data, the greater the collective’s ability to influence the associated label. The second term, here $-\frac{N-n}{N}(\Delta_{\tilde{x}}^{(n)}+2R_{\tilde{\delta}}(n)+2R_{\tilde{\delta}}(N-n))$, captures the counteracting influence of non-collective individuals in the population, quantifying how much they might limit the collective’s success in planting the signal. It indicates how strongly the target label is associated with the signal set: the more frequent $y^{*}$ is in the signal set, the easier it is to plant the signal; if other labels are far more likely than $y^{*}$, planting the signal becomes more difficult. The third term, $-\frac{\varepsilon}{1-\varepsilon}$, represents the platform’s ability to adapt to the distribution of its users. As $\varepsilon\mapsto\frac{\varepsilon}{1-\varepsilon}$ increases with $\varepsilon$, it benefits the collective to have $\varepsilon$ close to zero, limiting the platform’s flexibility and resulting in a tighter bound.
@@ 3.2.2 Feature-only signal planting
$\begin{split}\hat{S}(n)&\geq\underset{x^{\prime}\sim D^{(n)}}{\hat{\mathbb{P}}}\left[\frac{n}{N}\left(\underset{\tilde{D}^{(n)}}{\hat{\mathbb{P}}}(g(x^{\prime}),y^{*})-2R_{\tilde{\delta}}(n)\right)-\frac{N-n}{N}\left(\Delta_{g(x^{\prime})}^{(n)}+2R_{\tilde{\delta}}(n)+2R_{\tilde{\delta}}(N-n)\right)-\frac{\varepsilon}{1-\varepsilon}>0\right]\\ &\quad-R_{\tilde{\delta}}(n)-R_{\tilde{\delta}}(N_{\rm test}),\end{split}$ (2)
@@ 3.3.2 Adaptive strategy
The collective can also use an adaptive strategy, meaning that each member of the collective $(x,y)$ can change its feature-label pair to some $(g(x),y_{g(x)})$ where the modified label $y_{g(x)}$ depends on $g(x)$ and is no longer fixed. The plan for the collective is to estimate the optimal label using a subset of $n_{\rm est}<n$ randomly chosen participants. Formally, we assume that $D^{(n)}$ is the concatenation of two independent datasets $D^{(n_{\rm est})}$ and $D^{(n-n_{\rm est})}$ drawn from $\mathcal{D}$. A natural strategy for the collective is to change every pair $(x,y)$ into $(g(x),\hat{y}_{g(x)})$ where for $\tilde{x}\in\tilde{\mathcal{X}}$:
$\hat{y}_{\tilde{x}}:=\underset{y^{\prime}\in\mathcal{Y}\backslash\{y^{*}\}}{\text{argmax }}\underset{D^{(n_{\rm est})}}{\hat{\mathbb{P}}}(\tilde{x},y^{\prime}).$ (3)
@@
Let $\delta>0$, and write $\tilde{\delta}:=\delta/(2+6\#\tilde{\mathcal{X}})$. Let $n_{\rm est}<n$. Then, by playing the signal unplanting strategy above against a classifier that is $\varepsilon$-suboptimal on $\tilde{\mathcal{X}}$, the collective achieves with probability at least $1-\delta$ (over the draw of the consumers):
$\begin{split}\hat{S}(n)&\geq\underset{\tilde{x}\sim\tilde{D}^{(n)}}{\hat{\mathbb{P}}}\left[\frac{n}{N}\left(\underset{\tilde{D}^{(n)}}{\hat{\mathbb{P}}}(\tilde{x})-2R_{\tilde{\delta}}(n)\right)-\frac{N-n}{N}\left(\Delta_{\tilde{x}}^{(n-n_{\rm est})}+2R_{\tilde{\delta}}(n-n_{\rm est})+2R_{\tilde{\delta}}(N-n)\right)-\frac{\varepsilon}{1-\varepsilon}>0\right]\\ &\quad-R_{\tilde{\delta}}(n)-R_{\tilde{\delta}}(N_{\rm test}),\end{split}$ (4)
where $\Delta_{\tilde{x}}^{(n-n_{\rm est})}:=\underset{D^{(n-n_{\rm est})}}{\hat{\mathbb{P}}}(\tilde{x},y^{*})-\underset{D^{(n-n_{\rm est})}}{\hat{\mathbb{P}}}(\tilde{x},\hat{y}_{\tilde{x}})$.
@@ 3.4 Signal erasing
To understand Section 3.4, consider data poisoning in image classification. Section 3.4 holds in data poisoning strategies where the trigger is a fixed, opaque watermark, as in the case of binary masks Gu et al. (2019), where applying the mask twice is equivalent to applying it once. However, this is not true for strategies like pixel blending Chen et al. (2017). In contrast, Section 3.4 naturally applies to tabular data, where the collective can apply a transformation $g$ to map a feature to a constant value.
@@
$\begin{split}\hat{S}(n)&\geq\underset{x^{\prime}\sim D^{(n)}}{\hat{\mathbb{P}}}\left[\frac{n}{N}\left(\underset{D^{(n)}}{\hat{\mathbb{P}}}(x^{\prime})-2R_{\tilde{\delta}}(n)\right)-\frac{N-n}{N}\left(\Delta_{x^{\prime}}^{(n)}+2R_{\tilde{\delta}}(n)+2R_{\tilde{\delta}}(N-n)\right)-\frac{\varepsilon}{1-\varepsilon}>0\right]\\ &\quad-R_{\tilde{\delta}}(n)-R_{\tilde{\delta}}(N_{\rm test}),\end{split}$ (5)
@@ 4.1 Signal planting
Figure 2 shows how the lower bound varies with $N$. For a fixed fraction $n/N$, larger $N$ improves the collectives’s estimation ability as $n$ grows, reducing error terms $R_{\tilde{\delta}}(n)$. This suggests that platforms interacting with large user bases are more exposed to collectives altering their data. While this observation is specific to signal planting, it is even more relevant when optimal strategies need to be estimated, as in signal unplanting.
@@ 4.2 Signal unplanting
The lower bound obtained in Theorem 3.7 depends on $n_{\rm est}$, the size of the sub-collective used to determine the strategy $h$. As shown in Figure 3 (a), $n_{\rm est}$ involves a trade-off: small values lead to erratic strategy estimates and weaker bounds, while overly large values increase the $R_{\tilde{\delta}}(n-n_{\rm est})$ term, also weakening the bound. A good balance is achieved at $n_{\rm est}=2,000$.
[…]
In Figure 3 (c), we compare the lower bound with $n_{\rm est}=2,000$ and the actual success $\hat{S}(n)$ achieved at test time. The lower bound suggests that the collective would need to represent around $10\%$ of the agents interacting with the platform to have a significant impact, while in practice, only about $3\%$ was sufficient.
@@ Acknowledgements
The authors would like to thank Eugène Berta, Sacha Braun, Yurong Chen, and David Holzmüller for their helpful comments and suggestions. Funded by the European Union (ERC-2022-SYG-OCEAN-101071601). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This publication is part of the Chair “Markets and Learning”, supported by Air Liquide, BNP PARIBAS ASSET MANAGEMENT Europe, EDF, Orange and SNCF, sponsors of the Inria Foundation. This work has also received support from the French government, managed by the National Research Agency, under the France 2030 program with the reference “PR[AI]RIE-PSAI” (ANR-23-IACL-0008).
@@ Appendix A Additional Details on Experiments
For further details on the dataset composition, we refer to the code available at: https://github.com/GauthierE/statistical-collusion.
@@ Appendix B Algorithms
1:  Input: $\mathcal{X},\mathcal{Y},N,N_{\rm test},n<N,n_{\rm est}<n,D^{(n_{\rm est})},D^{(n-n_{\rm est})},g,y^{*},\delta>0,\varepsilon>0$
@@
3:  Observe $D^{(n_{\rm est})}$ and compute $\hat{y}_{\tilde{x}}:=\underset{y^{\prime}\in\mathcal{Y}\backslash\{y^{*}\}}{\text{argmax }}\underset{D^{(n_{\rm est})}}{\hat{\mathbb{P}}}(\tilde{x},y^{\prime})$ for every $\tilde{x}\in\tilde{\mathcal{X}}$
4:  Observe $D^{(n-n_{\rm est})}$ and compute $\Delta_{\tilde{x}}^{(n-n_{\rm est})}$ for every $\tilde{x}\in\tilde{\mathcal{X}}$
@@
6:  Compute $\tilde{D}^{(n)}$ by applying $h$ to all samples in $D^{(n)}$ defined as the concatenation of $D^{(n_{\rm est})}$ and $D^{(n-n_{\rm est})}$
@@
9:  Compute $R_{\tilde{\delta}}(n)$, $R_{\tilde{\delta}}(n-n_{\rm est})$, $R_{\tilde{\delta}}(N-n)$, and $R_{\tilde{\delta}}(N_{\rm test})$
10:  Compute and return:$\underset{\tilde{x}\sim\tilde{D}^{(n)}}{\hat{\mathbb{P}}}\left[\frac{n}{N}\left(\underset{\tilde{D}^{(n)}}{\hat{\mathbb{P}}}(\tilde{x})-2R_{\tilde{\delta}}(n)\right)-\frac{N-n}{N}\left(\Delta_{\tilde{x}}^{(n-n_{\rm est})}+2R_{\tilde{\delta}}(n-n_{\rm est})+2R_{\tilde{\delta}}(N-n)\right)-\frac{\varepsilon}{1-\varepsilon}>0\right]-R_{\tilde{\delta}}(n)-R_{\tilde{\delta}}(N_{\rm test})$

### v2

@@ 1 Introduction
In a network, agents often have diverse motivations and multiple incentives. When the incentives of the interacting agents do not fully align with those of the designer of the learning system, the former may wish to influence the learning process. This becomes salient in the common case in which a learning system interacts with a large number of agents. Such agents may collaborate, forming a cartel, by pooling their data and devising a common strategy. Accordingly, even if the learning algorithm is robust to single-agent adversarial behavior, the collective may be able to exert a significant influence on the algorithm. This concept of collective action has its origins in economic theories (Olson, 1965) and has been more recently explored in the context of machine learning (Hardt et al., 2023). It is crucial to understand how such collectives can influence learning algorithms to better comprehend the dynamics within the network of agents and to deploy algorithms that are reliable and aligned with the majority interests of consumers. We will focus specifically on the case of a platform that deploys a learning algorithm and interacts with a population of consumers.
@@ 2 Related Work
What sets the concept of collective action apart conceptually is its treatment of the collective as a group of individuals, each representing a data point. This perspective also has economic, social, and political dimensions, as certain groups of individuals can unite and collaborate to influence decisions. Such ideas have been explored in areas of research at the intersection of machine learning and other fields (Vincent et al., 2019; Albert et al., 2020; Vincent et al., 2021; Albert et al., 2021; Creager & Zemel, 2021; Vincent & Hecht, 2021). See Appendix A of Hardt et al. (2023) for an in-depth analysis of the related work on collective action. In addition, Ben-Dov et al. (2024) highlight how the learning algorithm shapes the success of collective action.
An important contribution of Hardt et al. (2023) is to study data poisoning via the formalism of Bayes-optimal classification, which yields a conceptual inversion of the idea of strategic classification (Hardt et al., 2016). While strategic classification revolves around a firm’s ability to anticipate and respond to the actions of a single, strategic individual, collective action shifts the focus toward a scenario where individuals collectively anticipate and strategically respond to the optimizing behavior of the firm. This concept has also been explored by Zrnic et al. (2021). Unlike traditional strategic classification, which primarily considers the firm’s perspective, collective action highlights the role of workers and consumers on online platforms.
@@ 3.2.1 Feature-label signal planting
$\begin{split}\hat{S}(n)&\geq\underset{\tilde{x}\sim\tilde{D}^{(n)}}{\hat{\mathbb{P}}}\left[\frac{n}{N}\left(\underset{\tilde{D}^{(n)}}{\hat{\mathbb{P}}}(\tilde{x})-2R_{\tilde{\delta}}(n)\right)\right.\\ &\left.\quad-\frac{N-n}{N}\left(\Delta_{\tilde{x}}^{(n)}+2R_{\tilde{\delta}}(n)+2R_{\tilde{\delta}}(N-n)\right)\right.\\ &\left.\quad-\frac{\varepsilon}{1-\varepsilon}>0\right]-R_{\tilde{\delta}}(n)-R_{\tilde{\delta}}(N_{\rm test}),\end{split}$ (1)
@@
Proofs and additional remarks can be found in Appendix E. Note that lower bound (1) is fully computable by the collective as it depends only on datasets $D^{(n)}$ and $\tilde{D}^{(n)}$. $D^{(n)}$ is obtained through data pooling by the collective’s members, and $\tilde{D}^{(n)}$ is computed by applying strategy $h$ to the dataset $D^{(n)}$. Details on the algorithm for computing the lower bound are provided in Appendix B, and the code is available at: https://github.com/GauthierE/statistical-algorithmic-action.
The interpretation of lower bound (1) is that success increases step by step: each feature $\tilde{x}$ in the signal set $\tilde{\mathcal{X}}$ has a certain resistance to being planted. As the relative size of the collective $n/N$ gradually increases, features $\tilde{x}$ are cracked as their resistance breaks, in decreasing order of resistance. For each feature $\tilde{x}$, its resistance depends on three terms. The first one, here $\frac{n}{N}(\underset{\tilde{D}^{(n)}}{\hat{\mathbb{P}}}(\tilde{x})-2R_{\tilde{\delta}}(n))$, represents how prevalent the feature $\tilde{x}$ is in the modified data: the more frequently $\tilde{x}$ appears in the poisoned data, the greater the collective’s ability to influence the associated label. The second term, here $-\frac{N-n}{N}(\Delta_{\tilde{x}}^{(n)}+2R_{\tilde{\delta}}(n)+2R_{\tilde{\delta}}(N-n))$, captures the counteracting influence of non-collective individuals in the population, quantifying how much they might limit the collective’s success in planting the signal. It indicates how strongly the target label is associated with the signal set: the more frequent $y^{*}$ is in the signal set, the easier it is to plant the signal; if other labels are far more likely than $y^{*}$, planting the signal becomes more difficult. The third term, $-\frac{\varepsilon}{1-\varepsilon}$, represents the platform’s ability to adapt to the distribution of its users. As $\varepsilon\mapsto\frac{\varepsilon}{1-\varepsilon}$ increases with $\varepsilon$, it benefits the collective to have $\varepsilon$ close to zero, limiting the platform’s flexibility and resulting in a tighter bound.
@@ 3.2.2 Feature-only signal planting
$\begin{split}\hat{S}(n)&\geq\underset{x^{\prime}\sim D^{(n)}}{\hat{\mathbb{P}}}\left[\frac{n}{N}\left(\underset{\tilde{D}^{(n)}}{\hat{\mathbb{P}}}(g(x^{\prime}),y^{*})-2R_{\tilde{\delta}}(n)\right)\right.\\ &\left.-\frac{N-n}{N}\left(\Delta_{g(x^{\prime})}^{(n)}+2R_{\tilde{\delta}}(n)+2R_{\tilde{\delta}}(N-n)\right)\right.\\ &\left.-\frac{\varepsilon}{1-\varepsilon}>0\right]-R_{\tilde{\delta}}(n)-R_{\tilde{\delta}}(N_{\rm test}),\end{split}$ (2)
@@ 3.3.2 Adaptive strategy
The collective can also use an adaptive strategy, meaning that each member of the collective $(x,y)$ can change its feature-label pair to some $(g(x),y_{g(x)})$ where the modified label $y_{g(x)}$ depends on $g(x)$ and is no longer fixed. The plan for the collective is to estimate the optimal label using a subset of $n_{\rm e}<n$ randomly chosen participants. Formally, we assume that $D^{(n)}$ is the concatenation of two independent datasets $D^{(n_{\rm e})}$ and $D^{(n-n_{\rm e})}$ drawn from $\mathcal{D}$. A natural strategy for the collective is to change every pair $(x,y)$ into $(g(x),\hat{y}_{g(x)})$ where for $\tilde{x}\in\tilde{\mathcal{X}}$:
$\hat{y}_{\tilde{x}}:=\underset{y^{\prime}\in\mathcal{Y}\backslash\{y^{*}\}}{\text{argmax }}\underset{D^{(n_{\rm e})}}{\hat{\mathbb{P}}}(\tilde{x},y^{\prime}).$ (3)
@@
Let $\delta>0$, and write $\tilde{\delta}:=\delta/(2+6\#\tilde{\mathcal{X}})$. Let $n_{\rm e}<n$. Then, by playing the signal unplanting strategy above against a classifier that is $\varepsilon$-suboptimal on $\tilde{\mathcal{X}}$, the collective achieves with probability at least $1-\delta$ (over the draw of the consumers):
$\begin{split}\hat{S}(n)&\geq\underset{\tilde{x}\sim\tilde{D}^{(n)}}{\hat{\mathbb{P}}}\left[\frac{n}{N}\left(\underset{\tilde{D}^{(n)}}{\hat{\mathbb{P}}}(\tilde{x})-2R_{\tilde{\delta}}(n)\right)\right.\\ &\left.-\frac{N-n}{N}\left(\Delta_{\tilde{x}}^{(n-n_{\rm e})}+2R_{\tilde{\delta}}(n-n_{\rm e})+2R_{\tilde{\delta}}(N-n)\right)\right.\\ &\left.-\frac{\varepsilon}{1-\varepsilon}>0\right]-R_{\tilde{\delta}}(n)-R_{\tilde{\delta}}(N_{\rm test}),\end{split}$ (4)
where $\Delta_{\tilde{x}}^{(n-n_{\rm e})}:=\underset{D^{(n-n_{\rm e})}}{\hat{\mathbb{P}}}(\tilde{x},y^{*})-\underset{D^{(n-n_{\rm e})}}{\hat{\mathbb{P}}}(\tilde{x},\hat{y}_{\tilde{x}})$.
@@ 3.4 Signal erasing
To understand Section 3.4, consider data poisoning in image classification. Section 3.4 holds in data poisoning strategies where the trigger is a fixed, opaque watermark, as in the case of binary masks (Gu et al., 2019), where applying the mask twice is equivalent to applying it once. However, this is not true for strategies like pixel blending (Chen et al., 2017). In contrast, Section 3.4 naturally applies to tabular data, where the collective can apply a transformation $g$ to map a feature to a constant value.
@@
$\begin{split}\hat{S}(n)&\geq\underset{x^{\prime}\sim D^{(n)}}{\hat{\mathbb{P}}}\left[\frac{n}{N}\left(\underset{D^{(n)}}{\hat{\mathbb{P}}}(x^{\prime})-2R_{\tilde{\delta}}(n)\right)\right.\\ &\left.-\frac{N-n}{N}\left(\Delta_{x^{\prime}}^{(n)}+2R_{\tilde{\delta}}(n)+2R_{\tilde{\delta}}(N-n)\right)\right.\\ &-\left.\frac{\varepsilon}{1-\varepsilon}>0\right]-R_{\tilde{\delta}}(n)-R_{\tilde{\delta}}(N_{\rm test}),\end{split}$ (5)
@@ 4.1 Signal planting
[…]
Figure 2 shows how the lower bound evolves with the total number of individuals $N$. When the fraction $n/N$ is held fixed, increasing $N$ leads to a larger collective size $n$. This, in turn, improves the collective’s ability to estimate key quantities, as reflected by the decreasing error terms $R_{\tilde{\delta}}(n)$. This suggests that platforms interacting with large user bases are more exposed to collectives altering their data. While this observation is specific to signal planting, it is even more relevant when optimal strategies need to be estimated, as in signal unplanting.
@@ 4.2 Signal unplanting
The lower bound obtained in Theorem 3.7 depends on $n_{\rm e}$, the size of the sub-collective used to determine the strategy $h$. As shown in Figure 3 (a), $n_{\rm e}$ involves a trade-off: small values lead to erratic strategy estimates and weaker bounds, while overly large values increase the $R_{\tilde{\delta}}(n-n_{\rm e})$ term, also weakening the bound. A good balance is achieved at $n_{\rm e}=2,000$.
@@
In Figure 3 (c), we compare the lower bound with $n_{\rm e}=2,000$ and the actual success $\hat{S}(n)$ achieved at test time. The lower bound suggests that the collective would need to represent around $10\%$ of the agents interacting with the platform to have a significant impact, while in practice, only about $3\%$ was sufficient.
@@ Acknowledgements
The authors would like to thank Eugène Berta, Sacha Braun, Yurong Chen, and David Holzmüller for their helpful comments and suggestions.
Funded by the European Union (ERC-2022-SYG-OCEAN-101071601). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
This publication is part of the Chair “Markets and Learning”, supported by Air Liquide, BNP PARIBAS ASSET MANAGEMENT Europe, EDF, Orange and SNCF, sponsors of the Inria Foundation.
This work has also received support from the French government, managed by the National Research Agency, under the France 2030 program with the reference “PR[AI]RIE-PSAI” (ANR-23-IACL-0008).
Impact Statement
This paper provides a theoretical and algorithmic foundation for studying collective influence, shedding light on the computations and coordination strategies necessary for such actions. By formalizing these mechanisms, we contribute to a deeper understanding of the risks and opportunities associated with collective interventions. This knowledge is essential for designing more robust, transparent, and socially aligned learning systems that account for strategic behavior and ensure fairer, more predictable outcomes in algorithmic decision-making.
@@ Appendix A Additional Details on Experiments
For further details on the dataset composition, we refer to the code available at: https://github.com/GauthierE/statistical-algorithmic-action.
@@ Appendix B Algorithms
1:  Input: $\mathcal{X},\mathcal{Y},N,N_{\rm test},n<N,n_{\rm e}<n,D^{(n_{\rm e})},D^{(n-n_{\rm e})},g,y^{*},\delta>0,\varepsilon>0$
@@
3:  Observe $D^{(n_{\rm e})}$ and compute $\hat{y}_{\tilde{x}}:=\underset{y^{\prime}\in\mathcal{Y}\backslash\{y^{*}\}}{\text{argmax }}\underset{D^{(n_{\rm e})}}{\hat{\mathbb{P}}}(\tilde{x},y^{\prime})$ for every $\tilde{x}\in\tilde{\mathcal{X}}$
4:  Observe $D^{(n-n_{\rm e})}$ and compute $\Delta_{\tilde{x}}^{(n-n_{\rm e})}$ for every $\tilde{x}\in\tilde{\mathcal{X}}$
@@
6:  Compute $\tilde{D}^{(n)}$ by applying $h$ to all samples in $D^{(n)}$ defined as the concatenation of $D^{(n_{\rm e})}$ and $D^{(n-n_{\rm e})}$
@@
9:  Compute $R_{\tilde{\delta}}(n)$, $R_{\tilde{\delta}}(n-n_{\rm e})$, $R_{\tilde{\delta}}(N-n)$, and $R_{\tilde{\delta}}(N_{\rm test})$
10:  Compute and return:$\underset{\tilde{x}\sim\tilde{D}^{(n)}}{\hat{\mathbb{P}}}\left[\frac{n}{N}\left(\underset{\tilde{D}^{(n)}}{\hat{\mathbb{P}}}(\tilde{x})-2R_{\tilde{\delta}}(n)\right)-\frac{N-n}{N}\left(\Delta_{\tilde{x}}^{(n-n_{\rm e})}+2R_{\tilde{\delta}}(n-n_{\rm e})+2R_{\tilde{\delta}}(N-n)\right)-\frac{\varepsilon}{1-\varepsilon}>0\right]-R_{\tilde{\delta}}(n)-R_{\tilde{\delta}}(N_{\rm test})$